How do you make 25 mL of a 7.8% saline solution from two solutions of 5.6% saline and 18.8% saline? Round to the hundredths.
Could you please show step to step on how it was solved. Thanks!

Answer by jorel1380(3719)

x+y=25
5.6x+18.8y=25x7.8=195
5.6x+5.6y=140
13.2y=55
y=4.167
x=20.833
You need 20.833ml of 5.6% solution and 4.167ml of 18.8% solution..